The sum of two numbers is 112. The second is 7 more than 4 times the first. What are the two numbers?
Lane invested $14,000 part as 15% and part at 2%. If the total interst at at end of year was $1060 how much did she invest at 15%?

Answer by checkley75(3666)

X+Y=112
X=112-Y
Y=4X+7 NOW SUBNSTITUTE (4X+7) FOR Y AND SOLVE FOR X IN THE FIRST EQUATION THUS:
X+(4X+7)=112
5X=112-7
5X=105
X=105/5
X=21 ANSWER FOR THE SMALLER DIGIT
Y=4*21+7
Y=84+7
Y=91
PROOF
91=4*21+7
91=84+7
91=91
AND 91+21=112